A water sample is not alkaline to phenolphthalein. However, 100 mL of the sample on titration with 0.02 N HCl required 16.0 mL to obtain end point using methyl orange as indicator. What are the types and amount of alkalinity present in the sample? 


1
Expert's answer
2021-08-25T06:36:30-0400

H2SO4 is dibasic acid so N\50 = 0.02N

Molarity of acid = 0.02\2 = 0.01M

So mol of acid = 10mL × 0.01M

= 0.1mmol

So amount of base = 2×0.1 mmol = 0.2mmol

Conc of base = (0.2×10-3 mol × 1000)\100

=2×10-3 M

POH = -log[OH-] = -log (2×10-3) =2.7

PH + POH = 14

PH = 14 -POH = 14 -2.7 = 11.3